# Neural Fields for Adaptive Photoacoustic Computed Tomography

### v2

@@ Abstract
Photoacoustic computed tomography (PACT) is a non-invasive imaging modality with wide medical applications. Conventional PACT image reconstruction algorithms suffer from wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue, which leads to image degradation. Accounting for these effects improves image quality, but measuring the SOS distribution is experimentally expensive. An alternative approach is to perform joint reconstruction of the initial pressure image and SOS using only the PA signals. Existing joint reconstruction methods come with limitations: high computational cost, inability to directly recover SOS, and reliance on inaccurate simplifying assumptions. Implicit neural representation, or neural fields, is an emerging technique in computer vision to learn an efficient and continuous representation of physical fields with a coordinate-based neural network. In this work, we introduce NF-APACT, an efficient self-supervised framework utilizing neural fields to estimate the SOS in service of an accurate and robust multi-channel deconvolution. Our method removes SOS aberrations an order of magnitude faster and more accurately than existing methods. We demonstrate the success of our method on a novel numerical phantom as well as an experimentally collected phantom and in vivo data. Our code and numerical phantom are available at https://github.com/Lukeli0425/NF-APACT.
Photoacoustic Computed Tomography, Inverse Problems, Neural Fields
@@ 1 Introduction
Photoacoustic computed tomography (PACT) images biological samples without cutting through tissue or disrupting normal function. Biological tissues absorb energy from laser stimulation and emit ultrasonic waves, which pass through the rest of the sample and are then collected by an external array of ultrasound transducers. An image of the initial pressure distribution (2D for the scope of this paper) is obtained by processing the ultrasonic signals collected by the transducer array. Traditional image reconstruction algorithms include filtered back-projection [1] and delay-and-sum (DAS) [2, 3, 4], which back-project the collected signals into the image using times of flight and assuming a constant SOS throughout the sample. This assumption neglects the impact of the sample’s internal structure and leads to image aberrations. Specifically, the inherent acoustic heterogeneity, or spatial variations in the 2D SOS distribution, will cause wavefront distortions in the ultrasonic waves that speed or delay their arrival to different parts of the transducer array. Visual artifacts of these distortions include ringing, smear, and doubling of features, which can obscure key sample structures.
A wide range of methods have been developed to mitigate or remove the artifacts caused by acoustic heterogeneity, as recently reviewed in [5, 6]. Autofocus approaches optimize an assumed-constant SOS based on the reconstructed image quality [7] or the coherence factor of the PA signals [8, 9]. Half-time reconstruction methods [10, 11] likewise assume a constant SOS but alleviate aberrations by identifying and discarding regions of the signals that are more affected by acoustic heterogeneity. Dual-SOS DAS [12] efficiently corrects for the first-order effect of acoustic heterogeneity by representing the sample as an oval-shaped region of altered SOS and assuming values for its shape and internally-constant SOS. On the other end of the complexity scale, Ultrasound Computed Tomography (USCT)-enhanced methods [13, 14, 15] directly measure the 2D SOS with a complementary ultrasound measurement apparatus. This measurement enhances image reconstruction but is challenging and expensive to implement simultaneously with PACT and USCT measurement procedures at the same time. The methods mentioned above either require oversimple assumptions or costly measurement of the SOS, and can only unreliably mitigate instead of fully undoing the SOS aberration in practice, see [5].
In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals. Regularization-based joint reconstruction [17, 18] alternatively reconstructs the image and SOS in an iterative framework with regularization on both distributions. However, the ill-posed nature of the problem causes numerical instability and error accumulation. The feature coupling method [19] parameterizes the SOS with several constant value regions and maximizes the similarity between reconstructed images from partial arrays. Its drawbacks are the inability to model a complicated SOS map and reliance on a good initial guess. Inspired by indirect wavefront sensing in adaptive optics, Adaptive PACT (APACT) [20] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search. APACT produces state-of-the-art image quality, but with two major drawbacks. First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image. Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application. Our joint reconstruction approach addresses these weaknesses with a novel deep learning approach.
Deep learning has been applied to SOS aberration correction [21, 22] and other image restoration tasks [23] in PACT. However, a key challenge lies in generalizing models trained on large simulated datasets to real-world in vivo data [6]. The scarcity of realistic, high-fidelity simulated data causes a domain shift between training and testing that can significantly degrade results on real data. We see a way around this challenge in the ability of neural networks to represent sample properties as functions of spatial coordinates, learned directly from observations in a self-supervised manner [16, 24]. These methods, termed implicit neural representations (INR) or neural fields (NF), have shown remarkable success in novel view synthesis [25, 26, 27], medical image reconstruction [28, 29, 30, 31, 32, 33], intensity diffraction tomography [34], and Fourier ptychographic microscopy [35]. NFs train a relatively small number of parameters with no need for external or labeled training data. This property is instrumental in scientific applications, where large labeled datasets are scarce.
Inspired by the success of NFs and APACT, we present Neural Fields for Adaptive Photoacoustic Computed Tomography (NF-APACT), a self-supervised framework for joint reconstruction of 2D SOS and initial pressure images in PACT, illustrated in Fig. 1. Our key contributions are:
@@
1. Our method achieves state-of-the-art image quality in a fraction of the compute time ( faster), while simultaneously recovering an approximate SOS. It is robust to nonideal effects (e.g. transducer EIRs and geometric offsets) found in real PACT systems, as demonstrated by its translation to in vivo data without modification.
2. We use no external training data, instead using only PA signals for self-supervised test-time training. Our method learns an SOS that reproduces the observed image aberrations by backpropagation through efficient and differentiable forward and inverse modules. The physics-based forward module incorporates a parameterization of the SOS by a neural field that reduces the number of unknowns by 2 orders of magnitude for fast and stable reconstruction.
3. NF-APACT is an interpretable method that explicitly reports the reconstructed SOS, applies a physics-based forward model from SOS to aberrations, and uses an analytical robust deconvolution for image reconstruction. Domain knowledge can be explicitly incorporated through user-defined regularization of the SOS.
We benchmark the proposed method and previous methods on image quality, compute time on both simulated and realistic data, showing that the proposed method performs better and faster compared to state-of-the-art method APACT [20]. Our code and numerical phantom are available in our GitHub repository (https://github.com/Lukeli0425/NF-APACT).
@@ 2.1 Image Formation and Aberration
If the SOS throughout the sample is constant and matches the background, there is a single velocity and delay that will create a sharp image. Because this is not the case for real samples, points in the object are smeared across the image into irregular patterns [36]. The delay value is a mitigation for these aberrations, and can improve image quality bending each back-projected arc, with the risk that poorly-selected delays will spread sample points into rings on the image (see PSFs in first column of Fig. 2). In real samples, with heterogeneous SOS, more complex aberrations arise and cannot be resolved with a single delay parameter (compare to right columns of Fig. 2). Specifically, assuming a straight acoustic rays model [36], the time-of-flight of a signal from sample point to a transducer at is
@@ 2.2 Multi-channel Deconvolution
Joint processing across multiple delay values provides more robust deconvolution results. A DAS reconstruction for each extra delay distance comprises our DAS stack, and for each patch of the image stack these delay values (in conjunction with SOS , on which more later) generates a corresponding PSF stack. Because each PSF takes the same underlying clean image into the aberrated image stack, the problem is overdetermined and can be expressed in the frequency domain as
Expressed in matrix form as , we reconstruct clean image with a pseudo-inverse,
@@
Fig. 3b shows the image patch recovered using single-channel deconvolution with each delay’s PSF. The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel. By using 32 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig. 3c. The reduction of aberrations as more delay channels are added can be seen in Fig. 3d, which also shows a growth in compute time from 0.4 to 1.1 minutes for 4 to 32 channels. While each multichannel deconvolution takes 5.0 seconds, this time is spent iteratively refining our estimate of the SOS, as described in the following sections.
2.3 Implicit Neural SOS Representation
To tackle the challenging problem of SOS reconstruction, we used a non-pretrained, coordinate-based, fully-connected network to parameterize the SOS as a function of sample location:
where are trained network parameters and is a set of spatial coordinates in the SOS map with the same resolution as the recovered image. To reduce computation, we restrict these coordinates to a circular mask roughly the size of the biological sample and set the SOS outside as a constant. This constant background SOS is calculated from the water temperature [37] and is also the SOS used in DAS.
The advantages of this approach are multi-fold. First, implicit neural representations can be learned in a self-supervised manner. This means the network is trained at test time for consistency on the input PA signals alone, which is vital due to the lack of datasets with ground truth images in this domain. Second, by learning a continuous function of the coordinates, the network serves as a natural prior that promotes continuity in the SOS map. Finally, the number of trained parameters is much smaller than the number of pixels to reconstruct in the SOS map. This significantly reduces the degrees of freedom in the inverse problem and makes the problem better-posed.
We used the same network architecture as SIREN [24], where sinusoidal activation functions are utilized to enhance the network’s ability to fit complex signals. Our MLP network consists of one hidden layer with 64 neurons and an output layer. The number of learnable parameters of our network is 4417, which is much smaller than the number of unknown values in the SOS map (200,000). Similar results were produced by networks with different neuron and layer counts, with 1.2-132k trainable parameters.
@@ 2.4 Joint Reconstruction
By combining our neural representation of SOS with differentiable PSF generation and deconvolution modules, shown in red in Fig. 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations. This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data. Specifically, we train the SOS network parameters to minimize the loss
over all image patches simultaneously. The first of these terms imposes a physics-informed data fidelity that includes both the forward model in as well as a Fourier-noise-specific weighting as in [20]. The second term smooths the SOS directly with total variation regularization [39]. Training is performed only on the SOS representation parameters but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered image patches . Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [20] for smooth interpatch transitions and adding the patches back to the initial pressure at corresponding patch locations.
We used Adam optimizer with a learning rate of and trained for 10 epochs. We select the learning rate and epoch count that empirically led to convergence in shortest compute time. Our method is implemented using PyTorch and is available at https://github.com/Lukeli0425/NF-APACT.
3 Experiments
@@ 3.1 Numerical Simulations
To test the proposed method with a ground truth reference, we designed a numerical phantom based on mouse liver structure, shown in Fig. 4a. The background water SOS is set to 1499.4 m/s (assuming a 26∘C water temperature based on [37]) outside the masked area, and SOS ranged from 1510.0 to 1650.0 m/s. The PA signals are obtained by 2D numerical simulation using the k-Wave toolbox [38] assuming an evenly distributed 512-transducer ring array with a diameter of 10 cm (same setting for the following phantom and in vivo experiments). Similar to [20], we take the derivative of the PA signals to simulate transducer effects: . The images are cropped into 75% overlapping 3.2mm 3.2mm patches, as for all reported results.
@@
The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [20], the wavefront errors can be decomposed into Fourier series, where the components degrade the image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (yellow), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (green). APACT includes only the zeroth and second order Fourier series terms in its wavefront model (blue), missing both the first order (the cosine component captured by dual-SOS in green) as well as higher order terms capturing finer detail. In comparison, NF-APACT can accurately model lower and higher-order components in the wavefront (purple), compared with the truth wavefront (red).
@@
These wavefront errors are directly reflected in image reconstruction quality in Fig. 4c. Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box). This shrinking is caused by the missing first order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue. Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, which significantly increases the reconstruction time in practice. Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported reconstruction time) and is able to fix the shrinking problem but artifacts still remain. The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term. APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts. Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods. Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores. Our reconstruction takes only 1.1 minutes, compared to 37 minutes for APACT. We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction.

### v3

# Self-Supervised Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography

@@ Abstract
Photoacoustic computed tomography (PACT) is a non-invasive imaging modality, similar to ultrasound, with wide-ranging medical applications. Conventional PACT images are degraded by wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue. Accounting for these effects can improve image quality and provide medically useful information, but measuring the SOS directly is burdensome and the existing joint reconstruction method is computationally expensive. Traditional supervised learning techniques are currently inaccessible in this data-starved domain. In this work, we introduce an efficient, self-supervised joint reconstruction method that recovers SOS and high-quality images using a differentiable physics model to solve the semi-blind inverse problem. The SOS, parametrized by either a pixel grid or a neural field (NF), is updated directly by backpropagation. Our method removes SOS aberrations more accurately and 35x faster than the current SOTA. We demonstrate the success of our method quantitatively in simulation and qualitatively on experimentally-collected and in-vivo data.
@@ 1 Introduction
Photoacoustic computed tomography (PACT) images biological samples without cutting through tissue or disrupting normal function [36]. Biological tissues absorb energy from laser stimulation and emit ultrasonic waves, which propagate through the rest of the sample and are then collected by an external array of ultrasound transducers. An image of the initial pressure (IP) distribution (2D for the scope of this paper) is obtained by processing the ultrasonic signals collected by the transducer array. Traditional image reconstruction algorithms include filtered back-projection [40] and delay-and-sum (DAS) [31, 10, 12], which back-project the collected signals into the image using times of flight assuming a constant SOS throughout the sample. This assumption neglects the impact of the sample’s internal structure and leads to image aberrations. Specifically, the inherent acoustic heterogeneity, or spatial variations in the 2D SOS distribution, will cause wavefront distortions in the ultrasonic waves that speed or delay their arrival to different parts of the transducer array. Visual artifacts of these distortions include ringing, smear, and doubling of features, which can obscure key sample structures [41].
Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available. We address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image. Specifically, as shown in in Fig. 1, we optimize a differentiable representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution. By imposing physical consistency across image locations and a delay-based “focal stack” (details in Sec. 3), this method converges quickly to produce an approximate SOS and high quality image. Our key contributions are:
1. A physics-based and interpretable PACT imaging model that is fully differentiable, enabling time- and memory-efficient optimization through backpropagation.
2. A more accurate, globally consistent representation of wavefront errors, for SOTA reconstruction quality in both image and SOS.
3. A self-supervised, test-time training framework that requires no external training data and can flexibly incorporate novel SOS representations and regularizers.
@@
We benchmark the proposed method and previous methods on simulated data, showing that our method achieves state-of-the-art reconstructions in a fraction of the compute time (40 sec vs. 23 min). In qualitative results on real data, we show the method is robust to nonideal effects (e.g. transducer responses and geometric offsets) found in PACT systems, as demonstrated by its translation to in vivo data without modification.
2 Related Work
Direct SOS Aberration Mitigation
A wide range of methods have been developed to mitigate or remove the artifacts caused by acoustic heterogeneity in PACT, as recently reviewed in [34, 30]. Autofocus approaches optimize an assumed-constant SOS based on the reconstructed image quality [33] or the coherence factor of the photoacoustic (PA) signals [43, 5]. Half-time reconstruction methods [1, 23] likewise assume a constant SOS but alleviate aberrations by identifying and discarding regions of the signals that are more affected by acoustic heterogeneity. Among these methods, Dual-SOS DAS [18] provides the best image reconstruction quality, by efficiently correcting the first-order effect of acoustic heterogeneity by representing the sample with an oval-shaped region of altered SOS and assuming values for its shape and internally-constant SOS. On the other end of the complexity scale, Ultrasound Computed Tomography (USCT)-enhanced methods [15, 19, 35] directly measure the 2D SOS with a complementary ultrasound measurement apparatus. This measurement enhances image reconstruction but is challenging and expensive to implement simultaneously with PACT. The above methods either require oversimple assumptions or costly measurement of the SOS, and can only unreliably mitigate instead of fully undoing the SOS aberration in practice [34].
Joint Recovery of SOS and Image
In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals. Regularization-based joint reconstruction [45, 25] alternatively reconstructs the image and SOS in an iterative framework with regularization on both distributions. However, the ill-posed nature of the problem causes numerical instability and error accumulation. The feature coupling method [3] parameterizes the SOS with several constant value regions and maximizes the similarity between reconstructed images from partial arrays. Its drawbacks are the inability to model a complicated SOS map and reliance on a good initial guess. Inspired by indirect wavefront sensing in adaptive optics (AO) [14, 9, 46], Adaptive PACT (APACT) [6] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search. Of these methods, APACT produces state-of-the-art image quality, but with two major drawbacks. First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image. Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application. Our method addresses these weaknesses with a novel self-supervised, physics-based deep learning approach.
Supervised Learning
Deep learning has been applied to SOS aberration correction [11, 13] and other image restoration tasks [7] in PACT. However, a key challenge lies in generalizing models trained on large simulated datasets to real-world in vivo data [30]. The scarcity of realistic, high-fidelity simulated data causes a domain shift between training and testing that can significantly degrade results on real data. Our physics-based framework, in contrast, shows robust performance across both simulated and in vivo data.
Neural Fields
Neural fields (NFs), or implicit neural representations (INRs) have been shown to be capable of representing a variety of natural signals [27, 38], and have been demonstrated in many computational imaging tasks, including medical imaging [21, 24, 44, 29, 26, 39]. NFs represent images with a relatively small number of parameters compared to pixel grids [28], with no need for external or labeled training data. Recent works in PACT have utilized NFs for IP image reconstruction from limited or sparse measurements [47, 37, 42], but these assume constant SOS and only reconstruct the IP image. NFs have also been used to reconstruct SOS in ultrasound CT [2] and have shown promise in estimating wavefronts for computational image reconstruction [16] and closed-loop control systems for adaptive optics [8]. We, like [16], do not intervene in the measurement process, but computationally correct for wavefront-based aberration parameters estimated with a NF. We use the SIREN NF architecture, which is a fully connected network with sinusoidal activations [27].
@@ 3.1 Image Formation and Aberration
If the SOS throughout the sample is constant and matches the background, there is a single velocity and delay that will create a sharp image. Because this is not the case for real samples, points in the object are smeared across the image into irregular patterns [41]. The delay value is a mitigation for these aberrations, and can improve image quality bending each back-projected arc, with the risk that poorly-selected delays will spread sample points into rings on the image (see PSFs in first column of Fig. 2). In real samples, with heterogeneous SOS, more complex aberrations arise and cannot be resolved with a single delay parameter (compare to right columns of Fig. 2). Specifically, assuming a straight acoustic ray model [41], the time-of-flight of a signal from sample point to a transducer at is
@@ 3.2 Multi-channel Deconvolution
Joint processing across multiple delay values provides more robust deconvolution results. A DAS reconstruction for each extra delay distance comprises our DAS stack, and for each patch of this image stack these delay values (in conjunction with SOS , on which more later) generate a corresponding PSF stack. Because each PSF takes the same underlying clean image into the aberrated image stack, the problem is overdetermined and can be expressed in the frequency domain as
and in matrix form as . We reconstruct clean image patches with a pseudo-inverse,
@@
Fig. 3b shows the image patch recovered using single-channel deconvolution with each delay’s PSF. The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel. By using 16 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig. 3c. The reduction of aberrations as more delay channels are added can be seen in Fig. 3d, which also shows a growth in compute time from 0.9 to 5.0 seconds for 4 to 32 channels.
@@ 3.3 Joint Reconstruction
By combining our differentiable PSF generation and deconvolution modules, shown in red in Fig. 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations. This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data by leveraging the redundant information encoded in the DAS stack.
We benchmark two types of SOS representation: a pixel grid (PG) with the same resolution as the IP image, and a neural field (NF) representation. For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [27]. This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions. The total number of NF parameters (1027) is much smaller than the number of pixels (200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed. See Sec. S3.3 for ablations on network size. For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [20]
We train the SOS parameters to minimize a self-supervised aberrated-image-matching loss:
over all image patches simultaneously, where is the parameters in the PG or NF. This loss function imposes a physics-informed data fidelity that includes both the forward model in as well as a Fourier-noise-specific weighting as in [6]. We use TV loss to regularize the conventional pixel grid representation, while for the NF representation, we rely instead on the network’s implicit regularization and set . Training is performed only on the SOS representation parameters but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered IP image patches . Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [6] for smooth interpatch transitions and adding the patches back to the IP image at corresponding patch locations. We use a delay count of by default in this paper (see ablations in Sec. 4.4). See Sec. S1.4 for details on our image patches and stitching.
4 Results
@@ 4.1 Numerical Simulations
To quantitatively benchmark the performance of our method with ground truth references, we designed a set of numerical phantoms with 5 different SOS maps (see Sec. S1.1 for details), one of which is shown in Fig. 4a for visualization of the results.
@@
The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [6], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow). APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [6] and Sec. S2 for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail. In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).
@@
These wavefront errors are directly reflected in image reconstruction quality in Fig. 4c. Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box). This shrinking is caused by the missing first-order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue. Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, significantly increasing the reconstruction time in practice. Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported time) and is able to fix the shrinking problem but artifacts remain. The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term. APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts coefficients. Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods. Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores. Our method with NF takes only 40 seconds, compared to 22 minutes for APACT. Our method with PG produces decent IP images in similar compute time but suffers from stripe-like SOS artifacts that originate from the path-integral nature of the wavefront calculation (see Eq. 3) in the straight acoustic ray model. We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction. See Tab. 1 for quantitative evaluation of all the methods, averaged over five numerical phantoms. Supplementary videos show the convergence of the IP image and SOS for each experiment.
| Method |
Time | ||||||||
| DAS | 21.49 | 0.372 | - | - | 0.13 s | ||||||||
| Dual-SOS DAS [18] | 24.42 | 0.446 | - | - | 0.12 s | ||||||||
| APACT [6] | 21.49 | 0.434 | 17.74 | 0.908 | 23 min | ||||||||
| Ours (PG) | 25.05 | 0.514 | 21.26 | 0.903 | 113.3 s | ||||||||
| Ours (NF) | 25.08 | 0.519 | 22.29 | 0.931 | 40.3 s | ||||||||
| Ours (SOS oracle) | 25.61 | 0.537 | - | - | 2.6 s |